# Structure-reinforced Transformer for Dynamic Graph Representation Learning with Edge Temporal States

### v2

@@ III-B Recurrent Temporal Node Feature Extraction
where $\mathcal{F}_{g_{l}}$ denotes an $l$-layer SGT, which is constructed by stacking $l$ encoding layers on top of each other. More specifically, during the first time slice, the SGT is provided with an initial embedding matrix $\textbf{X}\in\Re^{|V|\times d}$ as its input. In this matrix, each row indexed by node id signifies the initial representation of that particular node.
The encoding layer details are illustrated in Figure 2. The upper left portion of Figure 2 presents a global semantic encoding module, which takes $\textbf{H}^{t-1}$ as its input. This module employs a standard self-attention mechanism as described by [28] to compute the semantic correlation between each pair of nodes. The computation process is defined as follows:
$\textbf{H}^{0}=\textbf{X}$ (8)
$\textbf{Q}^{t}=\textbf{H}^{t-1}\textbf{W}_{Q}^{t}$ (9)
$\textbf{K}^{t}=\textbf{H}^{t-1}\textbf{W}_{K}^{t}$ (10)
@@
$\textbf{A}^{t}=\frac{\textbf{Q}^{t}(\textbf{K}^{t})^{T}}{\sqrt{d_{K}}}$ (12)
where $\textbf{W}_{Q}^{t},\textbf{W}_{K}^{t}\in\Re^{d\times d^{\prime}}$, and $\textbf{W}_{V}^{t}\in\Re^{d\times d}$ denote the trainable projection matrices. Concurrently, $\textbf{Q}^{t}$, $\textbf{K}^{t}$, and $\textbf{V}^{t}$ represent the attention query, key, and value matrices, respectively. The regular attention matrix is denoted by $\textbf{A}^{t}\in\Re^{|V|\times|V|}$, where $a_{ij}^{t}\in\textbf{A}^{t}$ represents the semantic correlation strength between nodes $v_{i}$ and $v_{j}$. For the sake of clarity in representation, we assume $d^{\prime}=d$ and consider the single-head self-attention. It is worth noting that the extension to multi-head self-attention is straightforward.
Despite the global semantic correlation, the togological structure of node pairs contributes strongly to node representation learning [24, 23]. To integrate this structural information into the transformer architecture, we have designed a structure encoding module, as depicted in the bottom right section of Figure 2. The degree of a node is a significant metric in graph theory, acting as a measure of its importance in the network. Similarly, the shortest path length between two nodes can provide insightful information about their topological correlation. As such, we propose a method to capture the topological dependencies between pairs of nodes. For each pair of nodes, $<v_{i},v_{j}>$, we assemble a topological attribute vector, denoted as $\textbf{attr}_{s}$. This vector encompasses three main components: the out-degree of the source node ($sd_{out}$), the in-degree of the target node ($td_{in}$), and the length of the shortest path connecting the pair of nodes ($spd$). Mathematically, this can be represented as follows:
@@
Once computed, the topological attribute vector $\textbf{attr}_{s}$ undergoes a transformation via a linear layer and normalization process. The output of this process is a higher dimension vector, $\textbf{r}^{t}_{(ij)s}\in\Re^{d_{e}}$, which we interpret as the topological feature of the node pair. This higher-dimensional representation allows for a more comprehensive analysis of the underlying topological relationship between the nodes.
For the representation of temporal edge states, we establish a shortest path matrix, denoted as $\textbf{ATTR}_{(ij)p}^{t}\in\Re^{2\times spd}$, for each node pair $<v_{i},v_{j}>$:
@@
Here, the first row, $\textbf{attr}_{e}^{t}$, corresponds to the edge type present on the path $\mathcal{p}_{v_{i}\rightarrow v_{j}}^{t}$ at the given time slice $t$. Meanwhile, each dimension in the second row, $\textbf{attr}_{w}^{t}$, signifies the associated weight of the edge. Each edge type and weight are represented by a real-valued trainable embedding vector, achieved through two distinct embedding layers. This leads to the generation of two matrices: $\textbf{attr}_{e}^{t}\mapsto\textbf{R}_{e}^{t}\in\Re^{spd\times d_{e}}$ and $\textbf{attr}_{w}^{t}\mapsto\textbf{R}_{w}^{t}\in\Re^{spd\times d_{e}}$. Here, $\textbf{R}_{e}^{t}$ symbolizes the embedding matrix associated with edge types, while $\textbf{R}_{w}^{t}$ signifies the embedding matrix linked with edge weights.
Subsequently, these matrices undergo an element-wise product operation as follows:
@@
where $\odot$ represents the Hadamard product. The resultant matrix $\textbf{R}_{p}^{t}\in\Re^{spd\times d_{e}}$ serves as the feature matrix for $\mathcal{p}_{v_{i}\rightarrow v_{j}}^{t}$. Each row within this matrix corresponds to the representation of its respective edge.
In line with the conventional Transformer design [28], we incorporate a position encoding into each row of $\textbf{R}_{p}^{t}$ to pinpoint the relative location of each edge in the associated path. This position encoding is calculated as follows:
@@
where $pos$ signifies the position of an edge in the path and $i$ represents the column index of a row vector. Next, we employ a one-dimensional convolutional layer to lessen the dimensionality of $\textbf{R}_{p}^{t}$. This process results in the final path feature $\hat{\textbf{r}}_{(ij)p}^{t}\in\Re^{d_{e}}$, which encapsulates the edge’s temporal state along the shortest path. Finally, we combine the topology feature $\textbf{r}^{t}_{(ij)s}$ and the path feature $\hat{\textbf{r}}_{(ij)p}^{t}$ to form the graph structure feature $\textbf{r}_{ij}^{t}\in\Re^{2d_{e}}$ as follows:
@@
where $\oplus$ indicates the concatenation of two vectors.
To infuse the structural information of the graph into the transformer model, as depicted in the upper right section of Figure 2, we propose a structure-aware attention mechanism. This mechanism is designed to modulate the standard attention score, $a_{ij}^{t}\in\textbf{A}^{t}$, in light of the graph’s structural features $\textbf{r}_{ij}^{t}$. We conceptualize this modulation process as an affine transformation of $a_{ij}^{t}$, mathematically represented as follows:
@@
where $\lambda_{ij}^{t}$, $\sigma_{ij}^{t}$ are the scale and offset coefficient at time slice $t$, respectively; $\hat{a}_{ij}\in\hat{\textbf{A}}^{t}$ is the structure-reinforced attention for $v_{i}$ and $v_{j}$. $\hat{\textbf{H}}^{t}\in\Re^{|V|\times d}$ indicates the final output of the transformer encoding layer, which encodes the global semantic correlation, topological dependency and edge temporal states.
III-C Model Training
We incorporate dynamic link prediction as an auxiliary task to facilitate the training and optimizing of the model parameters. For a candidate edge $e_{ij}^{t+1}$, we compute its feature $\textbf{h}_{e_{ij}}^{t+1}\in\Re^{d}$ as the absolute difference of node features: $|\textbf{h}_{i}^{t}-\textbf{h}_{j}^{t}|$. This feature is then input into a fully-connected neural network layer, which predicts whether the edge is positive or negative. The prediction process can be mathematically expressed as:
@@ III-C Model Training
where $\textbf{W}_{o}$ and $\textbf{b}_{o}$ symbolize the trainable weights, and $\hat{p}_{ij}^{t+1}$ denotes the predicted probability of the edge being positive. Cross entropy is adopted as the loss function, given by:
@@
where $\Theta$ encompasses all the trainable parameters, and $\lambda$ is a hyperparameter that manages the $L2$ regularization strength, serving as a countermeasure against overfitting. The adjustment and optimization of these parameters are conducted using the mini-batch AdamW optimizer [29].
IV Experiments
IV-A Experimental Setup and Datasets
@@
TABLE II: Statistics of the four datasets.
@@
We perform comprehensive experiments on four real-world dynamic graphs, each with unique characteristics:
  • twi-Tennis [30]: Originates from Twitter’s mention graphs during major tennis tournaments in 2017. Nodes are Twitter accounts, and edges represent mentions. Snapshots highlight the most mentioned accounts, with the frequency of mentions encoded as node labels.
  • CollegeMsg [31]: Built from private messages on the University of California, Irvine’s online social network. Nodes symbolize users, and edges are defined by messages sent from one user to another at specific times. The dataset is segmented into 88 snapshots taken at intervals of 190,080 seconds.
  • cit-HepTh [32]: Based on the e-print arXiv, focusing on 7,577 high energy physics theory papers and their 51,315 citations. Nodes are papers, edges denote citations over time. The dataset is organized into 78 snapshots, with paper abstracts represented through word2vec embeddings [15].
  • sx-MathOF [33]: Depicts temporal interactions within the Math Overflow forum of Stack Exchange. An edge indicates a user responding to or commenting on another user’s question or answer. This dataset is divided into 79 snapshots, each covering a span of 30 days.
A summary of the statistics for these four datasets is provided in Table II.
IV-B Competing Methods
To assess the efficacy of our proposed RSGT, we compare it with ten state-of-the-art benchmarks. These can be broadly classified into three groups: methods utilized for static graphs [11, 12, 16], techniques applicable to discrete dynamic graphs [7, 13, 8], and approaches suitable for continuous dynamic graphs [9, 34, 35, 10]. We provide a brief description of each below:
  • Node2vec [12]: It extends DeepWalk by incorporating biased random walks.
  • GraphSAGE [16]: It is a generalized inductive framework built on GNN.
TABLE III: Performance comparisons of dynamic link prediction among competing baselines on twi-Tennis and ColledgeMsg.
@@ IV-B Competing Methods
DeepWalk[11] 61.96±1.27 61.05±2.83 61.51±2.62 61.27±2.81 66.54±5.36 67.57±5.81 68.22±5.95 67.86±5.86
Node2vec[12] 62.03±0.67 62.36±1.73 62.88±1.83 62.54±1.77 65.82±4.12 68.91±3.47 69.36±3.53 69.10±3.50
GraphSAGE[16] 62.76±1.76 62.02±1.63 62.50±1.78 62.26±1.66 58.91±3.67 60.23±4.15 60.57±4.30 60.45±4.22
EvolveGCN[7] 64.73±0.64 63.51±0.92 64.12±1.01 63.80±0.98 63.27±4.42 65.62±4.64 65.37±4.81 65.44±4.72
CoEvoSAGE[13] 67.39±2.51 67.71±2.03 68.24±2.17 67.95±2.12 67.72±3.17 69.81±3.12 70.20±3.25 70.02±3.64
ROLAND [8] 68.45±2.08 68.61±1.74 69.03±1.81 68.80±1.74 69.44±2.86 70.19±2.58 70.53±2.63 70.32±2.64
CTDNE[9] 58.14±2.67 58.33±2.04 58.91±2.19 58.61±2.12 62.55±3.67 65.34±2.20 65.82±2.36 65.56±2.34
TGAT[34] 69.01±1.54 69.03±1.41 69.56±1.56 69.24±1.47 70.35±2.54 71.01±2.35 71.43±2.46 71.20±2.37
CAW[35] 71.35±1.68 71.15±1.41 72.03±1.46 71.57±1.42 73.04±2.11 72.71±1.95 73.23±2.01 72.95±1.98
TREND[10] 74.02±1.78 73.83±1.61 75.46±1.63 74.63±1.66 74.55±1.95 74.33±1.91 75.91±2.04 75.64±2.09
@@
Gains 18.33% 18.02% 16.71% 17.31% 16.45% 16.15% 14.89% 14.62%
TABLE IV: Performance comparisons of dynamic link prediction among competing baselines on cit-HepTh and sx-MathOF.
@@
DeepWalk[11] 51.55±0.90 49.89±0.92 50.89±0.89 50.39±0.98 66.23±0.89 66.47±0.94 67.81±0.98 67.14±1.12
Node2vec[12] 65.68±1.90 65.47±1.78 66.79±1.98 66.13±2.15 64.76±1.08 64.58±0.27 65.88±0.69 65.23±0.67
GraphSAGE[16] 70.72±1.96 70.56±2.12 71.98±2.04 71.27±2.41 65.32±1.55 65.52±1.35 66.84±1.69 66.18±1.21
EvolveGCN[7] 61.57±1.53 61.80±1.52 63.04±1.44 62.42±1.54 68.35±0.68 68.33±0.58 69.71±0.39 69.02±0.35
CoEvoSAGE[13] 65.88±1.32 65.98±1.28 67.32±1.19 66.65±0.98 71.38±1.02 71.37±0.92 72.81±0.87 72.09±0.96
ROLAND [8] 70.57±1.51 70.61±1.44 72.03±1.28 71.32±1.52 73.35±0.98 72.31±0.82 73.77±0.69 73.04±0.85
CTDNE[9] 49.42±1.86 43.79±1.98 44.67±2.67 44.23±2.92 60.72±0.28 60.30±0.39 61.52±0.19 60.91±0.51
TGAT[34] 71.42±1.28 70.39±1.30 71.81±1.36 71.10±1.34 74.15±1.01 74.00±0.81 75.50±0.83 74.75±0.79
CAW[35] 72.75±1.19 71.78±1.21 73.22±1.28 72.50±1.23 77.95±0.77 76.89±0.76 78.45±0.65 77.67±0.72
TREND[10] 80.37±2.08 80.32±1.97 81.94±1.82 81.13±1.92 79.82±1.56 79.21±1.27 80.81±1.31 80.01±1.34
@@
Gains 8.50% 8.13% 6.77% 7.44% 10.14% 10.40% 9.33% 9.86%
IV-C Experiment Results and Analyses
Prediction task and parameter settings

### v3

@@ 3.2 Recurrent Temporal Node Feature Extraction
where $\textbf{H}^{t}\in\Re^{|V|\times d}$ is the node feature matrix at time slice $t$, $\mathcal{F}_{g_{l}}(\cdot)$ denotes an $l$-layer SGT that constructed by stacking $l$ encoding layers, $|V|$ denotes the maximum node count across all time slices, and $d$ represents the dimensionality of the temporal node representation. Eq.7 implements a residual connection, facilitating gradient flow and enabling the model to leverage both current and historical information.
For the initial time slice ($t=1$), the SGT uses an initial embedding matrix $\textbf{H}^{0}=\textbf{X}\in\Re^{|V|\times d}$, where each row represents the initial feature vector of a particular node. This initial embedding can be derived from various sources, including node ID embeddings, where each node is assigned a unique ID embedded into a $d$-dimensional space; node attribute embeddings, which encode associated attributes such as user profiles in a social network; and pre-trained embeddings, which leverage pre-existing node representations from other models or domain knowledge as a starting point.
Next we will detail the core operations of our SGT, as shown in Figure 2, including global semantic encoding, structure encoding, and temporal feature extraction.
Global Semantic Encoding
The global semantic encoding module, illustrated in the upper left portion of Figure 2, is designed to capture the overall semantic relationships between nodes across the entire graph, regardless of their topological distance. This global perspective allows the model to understand node similarities and interactions beyond immediate neighborhoods, which is particularly crucial in dynamic graphs where distant nodes may become relevant over time.
To achieve this global semantic understanding, we employ a standard self-attention mechanism [30] to compute semantic correlations between all node pairs. This approach allows each node to attend to every other node in the graph, effectively creating a fully connected semantic graph layered on top of the existing topological structure. The module takes $\textbf{H}^{t-1}$ as input and processes it as follows:
$\textbf{A}^{t}=\frac{\textbf{Q}^{t}(\textbf{K}^{t})^{T}}{\sqrt{d_{K}}}=\frac{{\textbf{H}^{t-1}\textbf{W}_{Q}^{t}}({\textbf{H}^{t-1}\textbf{W}_{K}^{t}})^{T}}{\sqrt{d_{K}}}$ (8)
@@
where $\textbf{W}_{Q}^{t},\textbf{W}_{K}^{t}\in\Re^{d\times d^{\prime}}$, and $\textbf{W}_{V}^{t}\in\Re^{d\times d}$ are trainable projection matrices. $\textbf{Q}^{t}$, $\textbf{K}^{t}$, and $\textbf{V}^{t}$ represent the query, key, and value matrices, respectively. The attention matrix $\textbf{A}^{t}\in\Re^{|V|\times|V|}$ captures the semantic correlation strength between nodes, where $a_{ij}^{t}\in\textbf{A}^{t}$ represents the correlation between nodes $v_{i}$ and $v_{j}$. For simplicity, we assume $d^{\prime}=d$ and consider single-head self-attention, though extension to multi-head attention is straightforward.
Structure Encoding
While global semantic encoding captures overall node relationships, the topological structure of the graph provides crucial information for node representation learning [24, 23]. Our structure encoding module, depicted in the bottom right of Figure 2, incorporates this topological information into the transformer architecture, leveraging both local and global structural features to enhance the model’s capacity for capturing complex graph dynamics.
The judicious selection of topological features is paramount in effectively representing the structural characteristics of nodes within a graph. Drawing from seminal works in network analysis [31, 32], we incorporate node degree centrality and path-based measures, which have been demonstrated to encapsulate critical aspects of network topology. Specifically, node degree centrality, encompassing both in-degree and out-degree, serves as a fundamental metric reflecting a node’s connectivity and potential influence within the network, while the shortest path length between nodes captures structural proximity and information flow efficiency, thereby providing a comprehensive representation of a node’s position and role within the graph structure. Thus, for each node pair $\langle v_{i},v_{j}\rangle$, we construct a topological attribute vector $\textbf{attr}_{s}$:
@@
where $sd_{out}$, $td_{in}$, and $spd$ denote the out-degree of the source node, the in-degree of the target node, and the shortest path length between the nodes, respectively. Then we employ a transformation mechanism that projects this attribute vector into a higher-dimensional space, followed by a normalization step to ensure stable training dynamics:
$\textbf{r}^{t}_{(ij)s}=\text{LayerNorm}(\textbf{W}_{s}\textbf{attr}_{(ij)s}^{t}+\textbf{b}_{s})$ (11)
where $\textbf{W}_{s}\in\Re^{d_{e}\times 3}$ and $\textbf{b}_{s}\in\Re^{d_{e}}$ are learnable parameters, and LayerNorm denotes layer normalization. The resulting representation $\textbf{r}^{t}_{(ij)s}\in\Re^{d_{e}}$ encapsulates the complex topological relationships between nodes, facilitating the model’s ability to discern intricate structural patterns within the dynamic graph.
To additionally represent edge temporal states, we establish a shortest path matrix $\textbf{ATTR}_{(ij)p}^{t}\in\Re^{2\times spd}$ for each node pair, since it not only encodes the topological relationship between nodes but also incorporates information about how edge types and weights change over time, which is essential for understanding the dynamic nature of the graph:
@@
where $\textbf{attr}_{e}^{t}$ represents the edge types along the path $\mathcal{p}_{v_{i}\rightarrow v_{j}}^{t}$ at time $t$, and $\textbf{attr}_{w}^{t}$ represents the corresponding edge weights. To transform these discrete attributes into continuous representations suitable for neural network processing, we employ embedding layers:
$\textbf{attr}_{e}^{t}\mapsto\textbf{R}_{e}^{t}\in\Re^{spd\times d_{e}}$ (13)
$\textbf{attr}_{w}^{t}\mapsto\textbf{R}_{w}^{t}\in\Re^{spd\times d_{e}}$ (14)
where $d_{e}$ is the embedding dimension. These embeddings allow us to learn rich, continuous representations of edge types and weights, facilitating more nuanced modeling of temporal dynamics.
To capture the varying intensities of different edge types in the graph’s temporal evolution, we integrate edge type features with their corresponding weights. This integration is achieved through the Hadamard product, which effectively modulates the impact of each edge type based on its strength:
@@
where $\odot$ denotes the Hadamard product.
Since the position of edges along a path in a graph structure can carry crucial semantic information, and to enable our model to differentiate between identical edge types occurring at different positions, we incorporate positional encoding following the conventional Transformer design [30] that allows the model to consider the relative positions of edges, enhancing its ability to capture complex structural patterns:
@@
where $pos$ is the position of an edge in the path and $i$ is the column index.
We then apply a one-dimensional convolutional layer to reduce the dimensionality of $\textbf{R}_{p}^{t}$, resulting in the final path feature $\hat{\textbf{r}}_{(ij)p}^{t}\in\Re^{d_{e}}$. The graph structure feature $\textbf{r}_{ij}^{t}\in\Re^{2d_{e}}$ is formed by concatenating the topology feature and the path feature:
@@
where $\oplus$ denotes vector concatenation.
Temporal Feature Extraction
To incorporate the structural information into the transformer model, we propose a structure-aware attention mechanism, as shown in the upper right section of Figure 2. This mechanism modulates the standard attention score $a_{ij}^{t}\in\textbf{A}^{t}$ based on the graph’s structural features $\textbf{r}_{ij}^{t}$:
@@
where $\lambda_{ij}^{t}$ and $\sigma_{ij}^{t}$ are scale and offset coefficients, respectively. $\hat{a}_{ij}\in\hat{\textbf{A}}^{t}$ is the structure-reinforced attention for nodes $v_{i}$ and $v_{j}$. Thus the final output, i.e. the temporal feature $\hat{\textbf{H}}^{t}\in\Re^{|V|\times d}$, encodes global semantic correlation, topological dependency, and edge temporal states.
3.3 Model Training
To facilitate the training and optimization of model parameters, we incorporate dynamic link prediction as an auxiliary task. For a candidate edge $e_{ij}^{t+1}$, we compute its feature $\textbf{h}_{e_{ij}}^{t+1}\in\Re^{d}$ as the absolute difference of node features: $|\textbf{h}_{i}^{t}-\textbf{h}_{j}^{t}|$, which captures the relative differences in node representations and is crucial for predicting potential connections. The edge feature is then input into a fully-connected neural network layer for binary classification:
@@ 3.3 Model Training
where $\textbf{W}_{o}$ and $\textbf{b}_{o}$ are trainable weights, and $\hat{p}_{ij}^{t+1}$ denotes the predicted probability of the edge being positive. We then adopt cross entropy as the loss function:
@@
where $\Theta$ encompasses all trainable parameters, and $\lambda$ is a hyperparameter controlling the $L2$ regularization strength to mitigate overfitting.
For parameter optimization, we employ the mini-batch AdamW optimizer [33], which, by virtue of its efficient handling of large-scale graph data and its capacity to adapt learning rates dynamically, proves particularly well-suited to our model, where diverse parameter types and the evolving nature of graphs necessitate an approach that can accommodate both structural complexity and temporal variability.
3.4 Complexity Analysis
The time complexity of RSGT can be analyzed by examining its three main components. First, the Edge Temporal State Modeling, which constructs the weighted multi-relation graph, has a complexity of $O(|E|)$, where $|E|$ is the number of edges. This includes both edge type assignment and weight updating processes. The second and core component, Recurrent Temporal Node Feature Extraction, involves more intricate computations. For a graph with $|V|$ nodes and $d$-dimensional features, the global semantic encoding requires $O(|V|^{2}d)$ operations for self-attention computation. The structure encoding process, which includes subgraph sampling, shortest path computation, topological attribute calculation, and feature transformation, has a complexity of $O(|V|k(\log k+d))$, where $k$ is the subgraph sampling size. Notably, the use of subgraph sampling significantly reduces the computational burden from $O(|V|^{2}\log|V|)$ to $O(|V|k\log k)$, as typically $k\ll|V|$. The structure-aware attention mechanism adds an additional $O(|V|^{2}d)$ complexity. Considering $l$ layers in the SGT, the complexity for a single time step becomes $O(l(|V|k(\log k+d)+|V|^{2}d))$. Finally, the model training for dynamic link prediction contributes $O(|E|d)$ to the overall complexity. Combining these components and considering $T$ time steps, the total time complexity of RSGT is
$O(Tl(|V|k(logk+d)+|V|^{2}d)+|E|d)$ (27)
It’s worth noting that while the worst-case complexity appears to be quadratic in the number of nodes due to the global semantic encoding, our use of subgraph sampling significantly reduces the computational burden in practice, especially for the structure encoding part. This optimization allows RSGT to efficiently process large-scale dynamic graphs. In real-world scenarios where typically $k\ll|V|$ and $d\ll|V|$, the effective complexity is often much lower than the theoretical upper bound, enabling RSGT to scale well with increasing graph sizes.
4 Experiments
@@
4.1 Datasets
Table 1: Statistics of the four datasets.
@@ 4.1 Datasets
To evaluate RSGT’s performance across diverse scenarios, we conducted experiments on four real-world dynamic graph datasets: twi-Tennis, CollegeMsg, cit-HepTh, and sx-MathOF. These datasets span various domains and exhibit distinct characteristics in terms of time scales, node and edge densities, and interaction types.
The twi-Tennis dataset [34] captures Twitter mentions during major tennis tournaments, with nodes representing accounts and edges denoting mentions. CollegeMsg [35] is derived from a university’s online social network, where edges represent private messages between users. The cit-HepTh dataset [36] is based on arXiv’s high energy physics papers, with nodes as papers and edges as citations. Lastly, sx-MathOF [37] represents interactions on the Math Overflow forum, with edges indicating responses or comments to questions.
Each dataset is divided into multiple snapshots, capturing the temporal evolution of the respective networks. Node features, where available, are represented through appropriate embeddings. Table 1 summarizes the key statistics of these datasets.
Table 2: Performance comparisons of dynamic link prediction among competing baselines on twi-Tennis and ColledgeMsg.
@@
DeepWalk 61.96±1.27 61.05±2.83 61.51±2.62 61.27±2.81 66.54±5.36 67.57±5.81 68.22±5.95 67.86±5.86
GraphSAGE 62.76±1.76 62.02±1.63 62.50±1.78 62.26±1.66 58.91±3.67 60.23±4.15 60.57±4.30 60.45±4.22
EvolveGCN 64.73±0.64 63.51±0.92 64.12±1.01 63.80±0.98 63.27±4.42 65.62±4.64 65.37±4.81 65.44±4.72
CoEvoSAGE 67.39±2.51 67.71±2.03 68.24±2.17 67.95±2.12 67.72±3.17 69.81±3.12 70.20±3.25 70.02±3.64
ROLAND 68.45±2.08 68.61±1.74 69.03±1.81 68.80±1.74 69.44±2.86 70.19±2.58 70.53±2.63 70.32±2.64
CTDNE 58.14±2.67 58.33±2.04 58.91±2.19 58.61±2.12 62.55±3.67 65.34±2.20 65.82±2.36 65.56±2.34
TGAT 69.01±1.54 69.03±1.41 69.56±1.56 69.24±1.47 70.35±2.54 71.01±2.35 71.43±2.46 71.20±2.37
CAW 71.35±1.68 71.15±1.41 72.03±1.46 71.57±1.42 73.04±2.11 72.71±1.95 73.23±2.01 72.95±1.98
TREND 74.02±1.78 73.83±1.61 75.46±1.63 74.63±1.66 74.55±1.95 74.33±1.91 75.91±2.04 75.64±2.09
DyGFormer 75.89±1.52 75.67±1.43 77.21±1.58 76.43±1.50 76.32±1.78 76.11±1.69 77.54±1.86 76.82±1.77
R-GSAGE 74.85±1.63 74.62±1.55 76.18±1.59 75.39±1.57 75.21±1.83 74.98±1.75 76.43±1.92 75.70±1.83
ConTIG 75.13±1.59 74.91±1.48 76.49±1.55 75.69±1.51 75.78±1.81 75.54±1.72 77.02±1.89 76.27±1.80
@@
Gains 15.42% 15.14% 14.03% 14.55% 13.74% 13.47% 12.47% 12.86%
Table 3: Performance comparisons of dynamic link prediction among competing baselines on cit-HepTh and sx-MathOF.
@@
DeepWalk 51.55±0.90 49.89±0.92 50.89±0.89 50.39±0.98 66.23±0.89 66.47±0.94 67.81±0.98 67.14±1.12
GraphSAGE 70.72±1.96 70.56±2.12 71.98±2.04 71.27±2.41 65.32±1.55 65.52±1.35 66.84±1.69 66.18±1.21
EvolveGCN 61.57±1.53 61.80±1.52 63.04±1.44 62.42±1.54 68.35±0.68 68.33±0.58 69.71±0.39 69.02±0.35
CoEvoSAGE 65.88±1.32 65.98±1.28 67.32±1.19 66.65±0.98 71.38±1.02 71.37±0.92 72.81±0.87 72.09±0.96
ROLAND 70.57±1.51 70.61±1.44 72.03±1.28 71.32±1.52 73.35±0.98 72.31±0.82 73.77±0.69 73.04±0.85
CTDNE 49.42±1.86 43.79±1.98 44.67±2.67 44.23±2.92 60.72±0.28 60.30±0.39 61.52±0.19 60.91±0.51
TGAT 71.42±1.28 70.39±1.30 71.81±1.36 71.10±1.34 74.15±1.01 74.00±0.81 75.50±0.83 74.75±0.79
CAW 72.75±1.19 71.78±1.21 73.22±1.28 72.50±1.23 77.95±0.77 76.89±0.76 78.45±0.65 77.67±0.72
TREND 80.37±2.08 80.32±1.97 81.94±1.82 81.13±1.92 79.82±1.56 79.21±1.27 80.81±1.31 80.01±1.34
DyGFormer 81.25±1.89 81.18±1.82 82.73±1.75 81.95±1.78 80.94±1.43 80.36±1.21 81.89±1.28 81.12±1.24
R-GSAGE 81.58±1.85 81.52±1.79 83.08±1.71 82.29±1.75 81.63±1.38 81.02±1.17 82.57±1.23 81.79±1.20
ConTIG 82.19±1.92 82.13±1.85 83.72±1.78 82.92±1.81 81.21±1.41 80.62±1.19 82.16±1.25 81.38±1.22
@@
Gains 6.10% 5.75% 4.50% 5.13% 7.69% 7.94% 7.00% 7.47%
4.2 Competing Methods
To assess the efficacy of our proposed RSGT, we compare it with twelve state-of-the-art benchmarks, encompassing both static and dynamic graph representation learning approaches:
  • 1.
  • 2.
    GraphSAGE [39]: It is a generalized inductive framework built on GNN.
  • 3.
  • 4.
  • 5.
  • 6.
  • 7.
  • 8.
  • 9.
  • 10.
    DyGFormer [41]: It employs a Transformer-based architecture with neighbor co-occurrence encoding and patching techniques to learn from nodes’ historical first-hop interactions.
  • 11.
    R-GSAGE [17]: It captures the continuous dynamic evolution of node embedding trajectories using ordinary differential equations and self-attention mechanisms.
  • 12.
    ConTIG [20]: It integrates a recurrent structure into GraphSAGE to jointly explore structural and temporal patterns while maintaining a lightweight architecture.
4.3 Experiment Results and Analyses
4.3.1 Prediction task and parameter settings
Following the experimental framework outlined in [10], we use dynamic link prediction [42] to evaluate the efficacy of the proposed RSGT model.
Given a dynamic graph $\mathcal{G}$, we divide the snapshots from various time slices into training and test datasets. The training dataset $\mathcal{D}^{tr}$ encompasses all snapshots captured prior to time $tr$. These snapshots are utilized for training the representation learning models. Conversely, the snapshots captured post time $tr$ are allocated to the test dataset $\mathcal{D}^{te}$. To comprehensively evaluate our model’s capability in learning both short-term and long-term historical information, as well as its performance in predicting links over varying time horizons, we employ different train-test split ratios for each dataset, as shown in Tabel 1. This diverse set of split ratios allows us to simulate various real-world scenarios, from rapidly evolving social networks to more stable citation networks, thereby providing a more thorough assessment of RSGT’s adaptability and predictive power across different temporal scales.